Mr. Akpalu went to a market to purchase 3 kg of sugar, 10 kg of wheat and 1 kg of salt. In a shop near to Mr. Akpalu’s residence, these commodities are priced at € 20, €10 and € 6 per kg whereas in the local market these commodities are priced at € 15, € 8 and € 6 per kg respectively. If the cost of traveling to local market is € 25, find the net savings of Mr. Akpalu, using matrix multiplication method.

1
Expert's answer
2021-03-07T16:59:19-0500

Denote by "C_1" and "C_2" vectors of prices in a shop and in the market respectively. Then


"C_1=\begin{pmatrix}
 20 \\
 10 \\
6
\end{pmatrix},C_2=\begin{pmatrix}
 15 \\
 8 \\
6
\end{pmatrix}"

Let "Q" denotes vector of amounts needed for Mr.Akpalu. Then


"Q=\begin{pmatrix}
 3 \\
 10 \\
1
\end{pmatrix}"

Then the total price he pays equals to "Q(C_i)"

Thus the difference between total prices in 2 shops equals to


"Q(C_1)-Q(C_2)=Q(C_1-C_2)="

"\begin{pmatrix}
 3 & 10 & 1
\end{pmatrix}\bigg(\begin{pmatrix}
 20 \\
 10 \\
6
\end{pmatrix}-\begin{pmatrix}
 15 \\
 8 \\
6
\end{pmatrix}\bigg)"

"\begin{pmatrix}
 3 & 10 & 1
\end{pmatrix}\begin{pmatrix}
 5 \\
 2 \\
0
\end{pmatrix}=3(5)+10(2)+1(0)=35"

Since travelling to local market costs € 25 net savings equals to


"\u20ac 35-\u20ac 25=\u20ac 10"

The net saving of Mr.Akpalu is € 10.